# BlockLLM: Memory-Efficient Adaptation of LLMs by Selecting and Optimizing the Right Coordinate Blocks

### v1

@@ A.6 Pre-training on Llama
Table 9: Hyperparameter Details for Pre-training LLaMA Models with BlockLLM on the C4 Dataset
A.7 Ablation on the hyperparameter $m$
We investigated the sensitivity of the model to the patience parameter $m$ in both fine-tuning and pre-training setups. These experiments were conducted using the GLUE benchmark and the LLaMA 2 model on the C4 dataset. Throughout the experiments, we fixed all parameters of the Adam optimizer and maintained a sparsity level of $s=0.5$ while varying $m$. The results are presented in Figure 7. Our observations indicate that in the fine-tuning setting, the model is relatively insensitive to variations in $m$. Specifically, setting $m=50$ or $m=1000$ did not result in significant performance differences. This finding aligns with the observations reported in [34], which suggest that gradients change more slowly. The gradual nature of gradient changes implies a correspondingly gradual variation in the optimal parameter set, thereby reducing sensitivity to changes in $m$. In contrast, in the pre-training setting, smaller values of $m$ lead to faster convergence. This behavior can be attributed to the presence of noisy gradients in the earlier iterations of pre-training. Consequently, a smaller $m$ helps maintain impactful parameter selection particularly in the initial phase of training, thereby facilitating faster convergence.

### v2

@@ A.7 Pretraining on Llama
Table 10: Hyperparameter Details for Pretraining LLaMA Models with BlockLLM on the C4 Dataset
A.8 Ablation on the hyperparameter $m$
We investigated the sensitivity of the model to the patience parameter $m$ in both fine-tuning and pre-training setups. These experiments were conducted using the GLUE benchmark and the LLaMA 2 model on the C4 dataset. Throughout the experiments, we fixed all parameters of the Adam optimizer and maintained a sparsity level of $s=0.5$ while varying $m$. The results are presented in Figure 9. Our observations indicate that in the fine-tuning setting, the model is relatively insensitive to variations in $m$. Specifically, setting $m=50$ or $m=1000$ did not result in significant performance differences. This finding aligns with the observations reported in Zhao et al. (2024), which suggest that gradients change more slowly. The gradual nature of gradient changes implies a correspondingly gradual variation in the optimal parameter set, thereby reducing sensitivity to changes in $m$. In contrast, in the pre-training setting, smaller values of $m$ lead to faster convergence. This behavior can be attributed to the presence of noisy gradients in the earlier iterations of pre-training. Consequently, a smaller $m$ helps maintain impactful parameter selection particularly in the initial phase of training, thereby facilitating faster convergence.